OpenStudy (anonymous):

Two perpendicular lines intersect at the origin. If the slope of the first line is -1/2 what is the equation of the second line?

OpenStudy (johnweldon1993):

Perpendicular means negative reciprocal slopes If the slope of the first line is \(\large -\frac{1}{2}\) The second lines slope will be \[\large -\frac{1}{-\frac{1}{2}} = 2\]